a factory buys a new machine for $11,850. if the machine depreciates by 15% per year determine the depreciation charge in year 4 ?

A: $1,091.61
Answer #1

here, initial machine price, (p = $11850) Sinice depreiation occurs. Dy = ? Dy = P(1-84 = 11850 (1-0.154 2 11850 (0.8584 2 11